Grade 11Physical Chemistry

When 5 gm mixture of NaHCO3 and Na2CO3 is heatedd 560 ml of CO2 was collected at NTP.Calculate percentage composition of this mixture

Sheetal kumari
9 Years agoGrade 11

2 Answers

Goutham
ApprovedApproved Tutor Answer9 Years ago
Na2CO3 is a strong base and will not decompose  on normal heating. However NaHCO3 decomposes according to the equation:
2NaHCO3 → Na2CO3 + CO2  + H2O
No of moles of CO2 = 0.560/22400 =  0.025 moles
No of moles of NaHCO3 =  2*0.025 = 0.05 moles = 0.05*84 = 4.2 g
Weight of Na2CO3 = 5 – 4.2 = 0.8g
With this we can calculate the percentage composition
Ravimohan Kumar
8 Years ago
Ŕeaction is 2NaHCO3 ----> Na2CO3 + CO2 + H2Ono of moles of CO2 = 0.560/22.4 = 0.025 molesno of moles of NaHCO3 = 2 × 0.025 =0.05 moleswt. of NaHCO3 = 0.05 × 84 = 4.2gmwt. of Na2CO3 = 5 - 4.2 = 0.8gmso, % NaHCO3 = (4.2/5)×100 = 84%% Na2CO3 = (0.8/5)×100 = 16%